40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 27 15:46:13 2013, Jenne, Update, Locking, PRMI progress PRMI_sb_27Mar2013.png
    Reply  Thu Mar 28 01:59:33 2013, Jenne, Update, Locking, PRMI progress - even better! 
       Reply  Thu Mar 28 04:21:40 2013, rana, Update, Locking, PRMI progress - even better! 
Message ID: 8360     Entry time: Wed Mar 27 15:46:13 2013     Reply to this: 8363
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI progress 

[Gabriele, Jenne]

We have achieved PRMI locks of the order ~5 seconds! Here is an example lock:

PRMI_sb_27Mar2013.png

This was actuating MICH on the ITMs (+1 for ITMY, -1 for ITMX in the output matrix), and PRCL on PRM (+1). 

PRCL gain was +1, MICH gain was -10.

PRCL signal was normalized with POP22I with a matrix value of 0.003 . (No normalization of MICH).

Both PRCL and MICH were triggered on POP22I with high thresh of 200, low of 50.  MICH and PRCL FM2 (integrators) were triggered on POP22I with thresh of 400 and low thresh of 50.  FMs 4 and 5 were on for both MICH and PRCL always.

We zoomed in on the MICH_OUT signal, and the instability looks like it is around 300 Hz.  We aren't sure what this is.  I think this is a similar frequency to an oscillation that Yuta saw, but I'll have to check the old elogs.

PRM and BS SUS_LSC_POS filter banks both have notches between 1280-1290Hz.  The ITMs do not have this "Vio2" filter.

This is one of the first locks with the new triggering of both the INPUT and OUTPUT of the control filter banks.  I modified the lsc model before lunch. 

To do:  Where is this 300Hz coming from, and what can we do about it?  Why are we losing lock?  It's not due to the oscillation - maybe too much afternoon seismic?  Steve says he went next door and the rock monster / river is on medium/high.

 

ELOG V3.1.3-